OpenStudy (anonymous):

how do you find the domain and range of 4sin x?

OpenStudy (anonymous):

\[-1\le \sin(x) \le 1-> -4 \le 4.\sin(x) \le 4\] domain is all reals

OpenStudy (anonymous):

Range[-4,4] Domain R

OpenStudy (anonymous):

Is the range the amplitude up and down?

OpenStudy (anonymous):

@RaphaelFilgueiras

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

yw